If a refractive index of wateris 4/3 and the angle of incidence is 30. Find the speed of light in water
1
Expert's answer
2020-05-18T10:10:57-0400

Solution. From the definition of the refractive index


"n=\frac{c}{v}"

where c=3x10^8 m/s is  speed of light in vacuum, v is speed of light in material (water).

Therefore get


"v=\frac{c}{n}=\frac{3\times 10^9}{\frac{4}{3}}=2.25\times 10^9\frac{m}{s}"

Answer.

"v=2.25\times 10^9\frac{m}{s}"